OpenStudy (anonymous):

Surface Area / Volume question. Part 1: I need to create and provide the dimensions for two similar figures. (Rectangular Prisms) (Dimensions created below) Part 2: I also need to provide what the similarity ratio of these figures along with the ratio of their surface area and volume are. Part 3: Either using the actual volumes or using the formula, I need to show that the volume ratio is true.

OpenStudy (anonymous):

My shape of choice: rectangle dimensions L=2 W=3 similar rectangle is L=10 W=15

OpenStudy (unklerhaukus):

i think you need a 3 dimensional shape

OpenStudy (anonymous):

So should I add height to the dimensions?

OpenStudy (unklerhaukus):

sure

OpenStudy (anonymous):

Alright, L=2 W=3 H=5 L=10 W=15 H=25

OpenStudy (anonymous):

And we'll make it a rectangular prism i suppose.

OpenStudy (unklerhaukus):



OpenStudy (anonymous):

^ precisely.

OpenStudy (unklerhaukus):

ok so what is total surface area of each

OpenStudy (unklerhaukus):

\[SA=2lb+2bh+2hl\]

OpenStudy (unklerhaukus):

and the volume is \[V=lbh\]

OpenStudy (anonymous):

Sorry just got back to my computer. Let me figure these out real quick.

OpenStudy (anonymous):

SA1= 62 SA2= 1550

OpenStudy (anonymous):

V1=30 V2=3750

OpenStudy (anonymous):

So how does this look-- *1part: rectangle 1- L=2, W=3, H=5 rectangle 2- L=10, W=15, H=25 *2part: The ratio of these figures in similarity was 1:5 The SA to SA ratio was 1:25 The Volume to Volume ratio was 1:125 *3part: Formula (Surface Area) SA=2lb+2bh+2hl Formula (Volume)V=lwh so The first rectangle is 12+30+20={62(SA)} Volume = 30 The second rectangle is 300+750+500={1550(SA)} Volume = 3750 The second rectangle has area 25x the first rectangle {the scale was 5, and the area scale is 5^2 (2x5=10, 3x5=15 etc.)}

OpenStudy (anonymous):

ill change the rectangle -> rectangular prism but that was my template

OpenStudy (unklerhaukus):

looking great

OpenStudy (anonymous):

Thanks for the help!

OpenStudy (unklerhaukus):

So the surface area of similar rectangular prisms has as ratio of the square of the similarity ratio. And the ratio of their volumes is the third power of the similarity ratio.